determine the length of the diagonal of a square if its area is 256 units
Answer by stanbon(75887)

determine the length of the diagonal of a square if its area is 256 units
-----
Area of a square = side^2
---
s^2 = 256
s = sqrt(256) = 16
----
diagonal = sqrt[16^2 + 16^2]
= sqrt(2*16^2]
= 4*sqrt(2) units
========================
Cheers,
Stan H.
========================